# Deep Learning for Continuous-Time Stochastic Control with Jumps

### v2

@@ Abstract
In this paper, we introduce a model-based deep-learning approach to solve finite-horizon continuous-time stochastic control problems with jumps. We iteratively train two neural networks: one to represent the optimal policy and the other to approximate the value function. Leveraging a continuous-time version of the dynamic programming principle, we derive two different training objectives based on the Hamilton–Jacobi–Bellman equation, ensuring that the networks capture the underlying stochastic dynamics. Empirical evaluations on different problems illustrate the accuracy and scalability of our approach, demonstrating its effectiveness in solving complex, high-dimensional stochastic control tasks. Code is available at https://github.com/jdupret97/Deep-Learning-for-CT-Stochastic-Control-with-Jumps.
@@ 1 Introduction
for an initial condition $x\in D$, an $n$-dimensional Brownian motion $W$ and a controlled random measure $N^{\alpha}$ on $E\times\mathbb{R}_{+}$, with $E=\mathbb{R}^{l}\setminus\{0\}$, and suitable functions $\beta\colon[0,T]\times D\times A\to\mathbb{R}^{d}$, $\sigma\colon[0,T]\times D\times A\to\mathbb{R}^{d\times n}$ and $\gamma\colon[0,T]\times D\times E\times A\to\mathbb{R}^{d}$. The functions $f\colon[0,T]\times D\times A\to\mathbb{R}$ and $F\colon D\to\mathbb{R}$ model the running and final rewards, respectively. We assume the controlled random measure is given by $N^{\alpha}(B\times[0,t])=\sum_{j=1}^{M^{\alpha}_{t}}1_{\left\{Z_{j}\in B\right\}}$ for measurable subsets $B\subseteq E$, where $M^{\alpha}$ is a Poisson process with a stochastic intensity of the form $\lambda(t,X^{\alpha}_{t-},\alpha_{t})$ and $Z_{1},Z_{2},\dots$ are i.i.d. $E$-valued random vectors such that, conditionally on $\alpha$, the random elements $W,M^{\alpha}$ and $Z_{1},Z_{2},\dots$ are independent. Our goal is to find an optimal control $\alpha^{*}$ and the corresponding value of problem (1). In view of the Markovian nature of the dynamics (2), we work with feedback controls of the form $\alpha_{t}=\alpha(t,X^{\alpha}_{t-})$ for measurable functions $\alpha\colon[0,T)\times D\to A$ and consider the value function $V:[0,T]\times D\to\mathbb{R}$ given by
$V(t,x)=\sup_{\alpha}\mathbb{E}\!\left[\int_{t}^{T}f(s,X^{\alpha}_{s},\alpha_{s})ds+F(X^{\alpha}_{T})\,\Big|\,X^{\alpha}_{t}=x\right].$ (3)
@@
  • It yields accurate results for high-dimensional continuous-time stochastic control problems in cases where the dynamics of the underlying stochastic processes are known.
@@
While methods like finite differences, finite elements and spectral methods work well for solving partial (integro-)differential equations P(I)DEs in low dimensions, they suffer from the curse of dimensionality, and as a consequence, become infeasible in high dimensions. Recently, different deep learning based approaches for solving high-dimensional PDEs have been proposed (Raissi et al., 2017, 2019; Han et al., 2017, 2018; Sirignano & Spiliopoulos, 2018; Berg & Nyström, 2018; Beck et al., 2021; Lu et al., 2021; Bruna et al., 2024). They can directly be used to solve continuous-time stochastic control problems that admit an explicit solution for the optimal control in terms of the value function since in this case, the expression for the optimal control can be plugged into the HJB equation, which then reduces to a parabolic PDE. On the other hand, if the optimal control is not available in closed form, it cannot be plugged into the HJB equation, but instead, has to be approximated numerically while at the same time solving a parabolic PDE. Such implicit optimal control problems can no longer be solved directly with one of the deep learning methods mentioned above but require a specifically designed iterative approximation procedure. For low-dimensional problems with non-explicit optimal controls, a standard approach from the reinforcement learning (RL) literature is to use generalized policy iteration (GPI), a class of iterative algorithms that simultaneously approximate the value function and optimal control (Jacka & Mijatović, 2017; Sutton & Barto, 2018). Particularly popular are actor-critic methods going back to Werbos (1992), which have a separate memory structure to represent the optimal control independently of the value function. However, classical GPI schemes become impractical in high dimensions as the PDE and optimization problem both have to be discretized and solved for every point in a finite grid. This raises the need of meshfree methods for solving implicit continuous-time stochastic control problems in high dimensions with continuous action space. Several local approaches based on a time-discretization of (2) have been explored by e.g. Han & E (2016); Nüsken & Richter (2021); Huré et al. (2021); Bachouch et al. (2022); Ji et al. (2022); Li et al. (2024); Domingo-Enrich et al. (2024a, b). Alternatively, this can also be achieved using popular (deep) RL techniques including policy gradient methods such as A2C/A3C (Mnih et al., 2016), PPO (Schulman et al., 2017) and TRPO (Schulman et al., 2015); Q-learning type algorithms with DQN (Mnih et al., 2013), C51 (Bellemare et al., 2017), see also Wang et al. (2020); Jia & Zhou (2023); Gao et al. (2024); Szpruch et al. (2024); and hybrid approaches with DDPG (Lillicrap et al., 2019) or SAC (Haarnoja et al., 2018). However, these model-free deep RL algorithms do not explicitly take into account the underlying dynamics (2) of the stochastic control problem (1) but instead, solely on sampling from the environment. As a result, they are less accurate in cases where the state dynamics are known (see Figure 3 below). Moreover, this curse of dimensionality inherent in solving high-dimensional PDEs is further exacerbated when dealing with the jump-diffusion (2), as the resulting HJB PIDE (4) requires the numerical evaluation of jump expectations for every space-time point sampled from the domain.
@@
In this paper, we introduce a deep model-based approach for stochastic control problems with jumps which takes the system dynamics (2) into account by leveraging results derived from the HJB equation (4). This removes the need to simulate the underlying jump-diffusion equation and, as a result, avoids discretization errors. Our approach combines GPI with PIDE solving techniques and approximates the value function and optimal control in an actor-critic fashion with two neural networks trained iteratively on sampled data from the space-time domain. It has the advantage that it provides a global approximation of the value function and optimal control available for all space-time points, which can be evaluated rapidly in online applications. We develop two related algorithms. The first one, GPI-PINN, approximates the value function by training a neural network to minimize the residuals of the HJB equation (4), following a physics-inspired neural network (PINN) approach (Raissi et al., 2017, 2019) while leveraging Proposition 3.1 below to avoid the direct computation of the gradient $\nabla_{x}V$ and Hessian $\nabla_{x}^{2}V$ in the Hamiltonian (5). GPI-PINN can thus be viewed as an extension of the method proposed by Duarte et al. (2024), adapted to a finite-horizon setup with time-dependence and a terminal condition in the HJB equation, leading to time-dependent optimal control strategies and value function. It works well in high dimensions for control problems without jumps in the underlying dynamics (2) $(\gamma=0)$, but becomes inefficient in the presence of jumps as it requires computing the jump-expectation $\mathbb{E}\,V(t,x+\gamma(t,x,Z_{1},a))$ at numerous sample points $(t,x)$ in every iteration of the algorithm. To address this, our second algorithm, GPI-CBU, relies on a continuous-time Bellman updating rule to approximate the value function, thereby circumventing the computation of gradients, Hessians and jump-expectations altogether. This makes it highly efficient for high-dimensional stochastic control problems with jumps, even when the control is not available in closed form.
@@ 3 GPI-PINN
$\psi^{\prime\prime}(0)=\partial_{t}v(t,x)$ $+\beta^{\top}\hskip-1.42262pt(t,x,a)\,\nabla_{x}v(t,x)+\frac{1}{2}\text{Tr}\left[\sigma\sigma^{\top}\hskip-1.42262pt(t,x,a)\nabla^{2}_{x}v(t,x)\right].$ (7)
@@
In Algorithm 1, we simplify notations using $\mathcal{H}(t,x,\theta,\phi):=\mathcal{H}(t,x,V_{\theta},\alpha_{\phi}(t,x)).$
@@
for points $(z_{j})_{j=1}^{J}$ in $E$ sampled from the distribution $\mathcal{Z}$ of $Z_{1}$. Then in every gradient step, we approximate $\mathscr{L}_{1}$ and $\mathscr{L}_{2}$ by
$\widehat{\mathscr{L}}_{1}(\theta,\phi^{(k)})=\frac{\xi_{1}}{M_{1}}\sum_{m=1}^{M_{1}}\Big(\widehat{\mathcal{H}}\big(t_{m},x_{m},\theta,\phi^{(k)}\big)\Big)^{2}+\frac{\xi_{2}}{M_{2}}\sum_{m=1}^{M_{2}}\left(V_{\theta}(T,y_{m})-F(y_{m})\right)^{2}$ (14)
@@
$\widehat{\mathscr{L}}_{2}(\theta^{(k+1)},\phi)=-\frac{1}{M_{1}}\sum_{m=1}^{M_{1}}\widehat{\mathcal{H}}\hskip-1.42262pt\left(t_{m},x_{m},\theta^{(k+1)},\phi\right),\$(15)
@@
GPI-PINN then yields global approximations $V_{\theta^{(k_{*})}}$ and $\alpha_{\phi^{(k_{*})}}$ of the value function and optimal control on the whole space-time domain $0,T]\times D$. By using Proposition 3.1, it avoids the computation of the gradients and Hessians appearing in the Hamiltonian. However, it still has two drawbacks that make it inefficient for high-dimensional control problems with jumps. First, it has to approximate the jump-expectations $\mathbb{E}\big[V_{\theta}(t_{m},x_{m}+\gamma(t_{m},x_{m},Z_{1},\alpha_{\phi}(t_{m},x_{m})))\big]$ for all sample points $(t_{m},x_{m})$, $m=1,\dots,M_{1}$, in each of the gradient steps (16)–(17) and for every epoch $k=0,1,\dots,k_{*}$, see Eq. (13). Secondly, since the Hamiltonian is already a second-order integro-differential operator, the gradient steps $\nabla_{\theta}\widehat{\mathcal{H}}$ in (16) require the computation of third order derivatives, which is numerically costly.
@@ 4 GPI-CBU
$G_{\zeta}(t,x,z,v,a)=v(t,x)+\zeta$ $\big[\partial_{t}v(t,x)+f(t,x,a)+\beta^{\top}\hskip-1.42262pt(t,x,a)\,\nabla_{x}v(t,x)$ (18)
@@
$i=0,...,N_{1}-1$, to obtain $\theta^{(k+1)}=\theta^{(k)}_{N_{1}}$. To update the control network $\alpha_{\phi}$, we next introduce the operator
@@
Like GPI-PINN, GPI-CBU leverages Proposition 3.1 to bypass the computation of the gradients $\nabla_{x}V_{\theta}$ and Hessians $\nabla^{2}_{x}V_{\theta}$. In addition, it avoids the costly computation of the jump-expectations $\mathbb{E}\big[V_{\theta}\big(t_{m},x_{m}+\gamma(t_{m},x_{m},Z_{1},\alpha_{\phi}(t_{m},x_{m}))\big)\big]$ at each sample point $(t_{m},x_{m})$, $m=1,\ldots,M_{1}$. Also, when updating the value network using (22), it does not have to compute third-order derivatives like (16) since only $\nabla_{\theta}V_{\theta}$ is needed. This relates to the value update rule (19) of GPI-CBU being recursive as $G_{\zeta}$ now depends explicitly on the value weights $\theta^{(k)}$ computed at the previous epoch, in contrast to the residual-based GPI-PINN. On the other hand, since GPI-PINN averages over different jumps in each update, it exhibits more stable convergence than GPI-CBU; see the numerical results in Section 5 below. The proportionality factors $\xi_{1},\xi_{2}$ and the scaling factor $\zeta$ are hyperparameters. $\xi_{1}$ and $\xi_{2}$ can be fine-tuned following Wang et al. (2022). While Proposition 4.1 holds for any scaling factor $\zeta\in\mathbb{R}$, its choice has an influence on the numerical performance of GPI-CBU. In the numerical experiments of this paper, we set $\zeta=1$ as it provides a good trade-off between convergence speed and accuracy of the improvements in GPI-CBU. On the other hand, negative scaling factors always failed to converge to the true solutions with exploding losses $\widehat{\mathscr{L}}_{1}^{(k)}$ and $\widehat{\mathscr{L}}_{2}^{(k)}$ after only a few epochs. Alternatively, one could consider an adaptive scaling factor $\zeta_{k}>0$ depending on the epoch $k$. More details on hyperparameter fine-tuning are given in Appendix B. Finally, we emphasize that the policy updating rule of GPI-CBU (24) is equivalent to that of GPI-PINN (17), except that it circumvents the computation of the jump-expectations in $\widehat{\cal H}$ as it is now based on the expectation-free operator $\mathcal{G}$.

### v3

@@ Abstract
In this paper, we introduce a model-based deep-learning approach to solve finite-horizon continuous-time stochastic control problems with jumps. We iteratively train two neural networks: one to represent the optimal policy and the other to approximate the value function. Leveraging a continuous-time version of the dynamic programming principle, we derive two different training objectives based on the Hamilton–Jacobi–Bellman equation, ensuring that the networks capture the underlying stochastic dynamics. Empirical evaluations on different problems illustrate the accuracy and scalability of our approach, demonstrating its effectiveness in solving complex high-dimensional stochastic control tasks. Code is available at https://github.com/jdupret97/Deep-Learning-for-CT-Stochastic-Control-with-Jumps.
@@ 1 Introduction
for an initial condition $x\in D$, an $n$-dimensional Brownian motion $W$ and a controlled random measure $N^{\alpha}$ on $E\times\mathbb{R}_{+}$, with $E=\mathbb{R}^{l}\setminus\{0\}$, and suitable functions $\beta\colon[0,T]\times D\times A\to\mathbb{R}^{d}$, $\sigma\colon[0,T]\times D\times A\to\mathbb{R}^{d\times n}$ and $\gamma\colon[0,T]\times D\times E\times A\to\mathbb{R}^{d}$. The functions $f\colon[0,T]\times D\times A\to\mathbb{R}$ and $F\colon D\to\mathbb{R}$ model the running and final rewards, respectively. We assume the controlled random measure is given by $N^{\alpha}(B\times[0,t])=\sum_{j=1}^{M^{\alpha}_{t}}1_{\left\{Z_{j}\in B\right\}}$ for measurable subsets $B\subseteq E$, where $M^{\alpha}$ is a Cox process with a stochastic intensity of the form $\lambda(t,X^{\alpha}_{t-},\alpha_{t})$ and $Z_{1},Z_{2},\dots$ are i.i.d. $E$-valued random vectors such that, conditionally on $\alpha$, the random elements $W,M^{\alpha}$ and $Z_{1},Z_{2},\dots$ are independent. Our goal is to find an optimal control $\alpha^{*}$ and the corresponding value of problem (1). In view of the Markovian nature of the dynamics (2), we work with feedback controls of the form $\alpha_{t}=\alpha(t,X^{\alpha}_{t-})$ for measurable functions $\alpha\colon[0,T)\times D\to A$ and consider the value function $V:[0,T]\times D\to\mathbb{R}$ given by
$V(t,x)=\sup_{\alpha}\mathbb{E}\!\left[\int_{t}^{T}f(s,X^{\alpha}_{s},\alpha_{s})ds+F(X^{\alpha}_{T})\,\Big|\,X^{\alpha}_{t}=x\right].$
@@
  • It yields accurate results for high-dimensional continuous-time stochastic control problems in cases where the underlying system dynamics are known.
@@
While methods like finite differences, finite elements and spectral methods work well for solving partial (integro-)differential equations P(I)DEs in low dimensions, they suffer from the curse of dimensionality and, as a consequence, become infeasible in high dimensions. Recently, different deep learning based approaches for solving high-dimensional PDEs have been proposed (Raissi et al., 2017, 2019; Han et al., 2017, 2018; Sirignano and Spiliopoulos, 2018; Berg and Nyström, 2018; Beck et al., 2021; Lu et al., 2021; Bruna et al., 2024). They can directly be used to solve continuous-time stochastic control problems that admit an explicit solution for the optimal control in terms of the value function since in this case, the expression for the optimal control can be plugged into the HJB equation, which then reduces to a parabolic PDE. On the other hand, if the optimal control is not available in closed form, it cannot be plugged into the HJB equation, but instead, has to be approximated numerically while at the same time solving a parabolic PDE. Such implicit optimal control problems can no longer be solved directly with one of the deep learning methods mentioned above but require a specifically designed iterative approximation procedure. For low-dimensional problems with non-explicit optimal controls, a standard approach from the reinforcement learning (RL) literature is to use generalized policy iteration (GPI), a class of iterative algorithms that simultaneously approximate the value function and optimal control (Jacka and Mijatović, 2017; Sutton and Barto, 2018). Particularly popular are actor-critic methods going back to Werbos (1992), which have a separate memory structure to represent the optimal control independently of the value function. However, classical GPI schemes become impractical in high dimensions as the PDE and optimization problem both have to be discretized and solved for every point in a finite grid. This raises the need for meshfree methods to solve implicit continuous-time stochastic control problems in high dimensions with continuous action space. Several local approaches based on a time-discretization of (2) have been explored by e.g. Han and E (2016), Nüsken and Richter (2021), Huré et al. (2021), Bachouch et al. (2022), Ji et al. (2022), Li et al. (2024), Domingo-Enrich et al. (2024a), Domingo-Enrich et al. (2024b). Alternatively, (deep) RL methods can be used, for instance, Q-learning type algorithms such as DQN (Mnih et al., 2013) or C51 (Bellemare et al., 2017), see also Wang et al. (2020), Jia and Zhou (2023), Gao et al. (2024), Szpruch et al. (2024); or actor critic approaches such as DDPG (Lillicrap et al., 2019), SAC (Haarnoja et al., 2018), A2C/A3C (Mnih et al., 2016), PPO (Schulman et al., 2017) or TRPO (Schulman et al., 2015). However, these RL algorithms are model-free, that is, they do not explicitly take into account the underlying dynamics (2) of the control problem (1) but instead, solely rely on sampling from the environment. As a result, they are less accurate in cases where the system dynamics are known (see Figure 3 below). On the other hand, the difficulty of solving high-dimensional PDEs is further exacerbated if the system dynamics includes stochastic jumps since in this case, the HJB equation (3) requires the computation of the jump expectations $\mathbb{E}\,V(t,x+\gamma(t,x,Z_{1},a))$ for every space-time point $(t,x)$ sampled from the domain.
@@
In this paper, we introduce a deep model-based approach for stochastic control problems with jumps that takes the system dynamics (2) into account by leveraging the HJB equation (3). This removes the need to simulate the underlying jump-diffusion (2) and, as a result, avoids discretization errors. Our approach combines GPI with a PIDE solving method. It approximates the value function and optimal control in an actor-critic fashion with two neural networks trained iteratively on sampled data from the space-time domain. As such, it has the advantage that it provides global approximations of the value function and optimal control available for all space-time points, which can be evaluated rapidly in online applications. We develop two related algorithms. The first one, GPI-PINN, approximates the value function by training a neural network to minimize the residuals of the HJB equation (3), following a physics-inspired neural network (PINN) approach (Raissi et al., 2017, 2019) while leveraging Proposition 3.1 below to avoid the direct computation of the gradient $\nabla_{x}V$ and Hessian $\nabla_{x}^{2}V$ in the Hamiltonian (4). GPI-PINN can be viewed as an extension of the method proposed by Duarte et al. (2024) adapted to a finite-horizon setup with time-dependence and a terminal condition in the HJB equation, leading to time-dependent optimal control strategies and value function, see also Dupret and Hainaut (2024). It works well in high dimensions for control problems without jumps in the underlying dynamics (2) $(\gamma=0)$, but becomes inefficient in the presence of jumps as it requires the computation of the jump-expectation $\mathbb{E}\,V(t,x+\gamma(t,x,Z_{1},a))$ at numerous sample points $(t,x)$ in every iteration of the algorithm. To address this, our second algorithm, GPI-CBU, relies on a continuous-time Bellman updating rule to approximate the value function, thereby circumventing the computation of gradients, Hessians and jump-expectations altogether. This makes it highly efficient for high-dimensional stochastic control problems with jumps, even when the control is not available in closed form.
@@ 3 GPI-PINN
$\psi^{\prime\prime}(0)=\partial_{t}v(t,x)$ $+\beta^{\top}\hskip-1.42262pt(t,x,a)\,\nabla_{x}v(t,x)+\frac{1}{2}\text{Tr}\left[\sigma\sigma^{\top}\hskip-1.42262pt(t,x,a)\nabla^{2}_{x}v(t,x)\right].$
@@
In Algorithm 1, we simplify the notation by using $\mathcal{H}(t,x,\theta,\phi):=\mathcal{H}(t,x,V_{\theta},\alpha_{\phi}(t,x)).$
@@
for points $(z_{j})_{j=1}^{J}$ in $E$ sampled from the distribution $\mathcal{Z}$ of $Z_{1}$. Then, we approximate $\mathscr{L}_{1}$ and $\mathscr{L}_{2}$ in every time step by
$\widehat{\mathscr{L}}_{1}(\theta,\phi^{(k)})=\frac{\xi_{1}}{M_{1}}\sum_{m=1}^{M_{1}}\Big(\widehat{\mathcal{H}}\big(t_{m},x_{m},\theta,\phi^{(k)}\big)\Big)^{2}+\frac{\xi_{2}}{M_{2}}\sum_{m=1}^{M_{2}}\left(V_{\theta}(T,y_{m})-F(y_{m})\right)^{2}$
@@
$\widehat{\mathscr{L}}_{2}(\theta^{(k+1)},\phi)=-\frac{1}{M_{1}}\sum_{m=1}^{M_{1}}\widehat{\mathcal{H}}\hskip-1.42262pt\left(t_{m},x_{m},\theta^{(k+1)},\phi\right),\$
@@
GPI-PINN yields global approximations $V_{\theta^{(k_{*})}}$ and $\alpha_{\phi^{(k_{*})}}$ of the value function and optimal control on the whole space-time domain $0,T]\times D$. By using Proposition 3.1, it avoids the computation of the gradients and Hessians appearing in the Hamiltonian (4). However, it still has two drawbacks that make it inefficient for high-dimensional control problems with jumps. First, it has to approximate the jump-expectations $\mathbb{E}\big[V_{\theta}(t_{m},x_{m}+\gamma(t_{m},x_{m},Z_{1},\alpha_{\phi}(t_{m},x_{m})))\big]$ for all sample points $(t_{m},x_{m})$, $m=1,\dots,M_{1}$, in each of the gradient steps (10)–(11) and all epochs $k=0,1,\dots,k_{*}$; see (9). Secondly, since the Hamiltonian is already a second-order integro-differential operator, the gradient steps $\nabla_{\theta}\widehat{\mathcal{H}}$ in (10) require the computation of third order derivatives, which is numerically costly.
@@ 4 GPI-CBU
$G_{\zeta}(t,x,z,v,a)=v(t,x)+\zeta$ $\big[\partial_{t}v(t,x)+f(t,x,a)+\beta^{\top}\hskip-1.42262pt(t,x,a)\,\nabla_{x}v(t,x)$
@@
$i=0,...,N_{1}-1$, to obtain $\theta^{(k+1)}=\theta^{(k)}_{N_{1}}$. To update the control network $\alpha_{\phi}$, we introduce the operator
@@
Like GPI-PINN, GPI-CBU leverages Proposition 3.1 to bypass the computation of the gradients $\nabla_{x}V_{\theta}$ and Hessians $\nabla^{2}_{x}V_{\theta}$. In addition, it avoids the costly computation of the jump-expectations (9) of $V_{\theta}$ at each sample point $(t_{m},x_{m})$. Also, it it does not have to compute third-order derivatives like (10) when updating the value network since only $\nabla_{\theta}V_{\theta}$ is needed in (15). This is a consequence of the fact that the value update rule (15) of GPI-CBU is recursive as $G_{\zeta}$ depends on the value weights $\theta^{(k)}$ computed in the previous epoch in contrast to the residual-based GPI-PINN. On the other hand, since GPI-PINN averages over different jumps in each update, it exhibits more stable convergence than GPI-CBU; see Figure 2(a) below. The proportionality factors $\xi_{1},\xi_{2}$ and the scaling factor $\zeta$ are hyperparameters. $\xi_{1}$ and $\xi_{2}$ can be fine-tuned following Wang et al. (2022). While Proposition 4.1 holds for any scaling factor $\zeta\in\mathbb{R}$, its choice affects the numerical performance of GPI-CBU. In the numerical experiments of this paper, we set $\zeta=1$ as it provides a good trade-off between convergence speed and accuracy of the improvements in GPI-CBU. On the other hand, negative scaling factors failed to converge to the true solution in all our experiments with exploding losses $\widehat{\mathscr{L}}_{1}^{(k)}$ and $\widehat{\mathscr{L}}_{2}^{(k)}$ after only a few epochs. Alternatively, one could consider an adaptive scaling factor $\zeta_{k}>0$ depending on the epoch $k$. More details on hyperparameter fine-tuning are given in Appendix B. Finally, we note that the policy updating rule of GPI-CBU (17) is equivalent to that of GPI-PINN (11), except that it circumvents the computation of the jump-expectations in $\widehat{\cal H}$ since it uses the expectation-free operator $\mathcal{G}$.